# Scale Where It Matters: Training-Free Localized Scaling for Diffusion Models

### v2

@@ G.4 Practical Q&A
A2:   As in Corollary 3, when local and global trials cost the same, the comparison becomes purely statistical. In this case, the dominance condition in Corollary 1 becomes
@@
A3:   As shown by the sparse-regime analysis following Corollary 3, the sparse-defect regime strongly favors LoTTS because global resampling risks harming every clean patch. If $s\ll M$, the global-harm term
@@
A4:   As in equation 22, LoTTS only requires that each selected patch provides positive expected net gain:
$\pi>\Big(1+\tfrac{q\overline{\delta}}{h_{\ell}\overline{\gamma}}\Big)^{-1}.$
@@
A5:   As in equation G.3.1, when $C_{\ell}\approx C_{g}$, LoTTS needs

### v3

@@ G.4 Practical Q&A
A2: As in Corollary 5, when local and global trials cost the same, the comparison becomes purely statistical. In this case, the dominance condition in Corollary 3 becomes:
@@
A3: As shown by the sparse-regime analysis following Corollary 5, the sparse-defect regime strongly favors LoTTS because global resampling risks harming every clean patch. If $s\ll M$, the global-harm term:
@@
A4: As in Eq. equation 23, LoTTS only requires that each selected patch provides positive expected net gain:
$\pi>\bigl(1+\tfrac{q\overline{\delta}}{h_{\ell}\overline{\gamma}}\bigr)^{\!-1}.$
@@
A5: As in Eq. equation G.3, when $C_{\ell}\approx C_{g}$, LoTTS needs: